OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

so u obviously have to distribute 1st

OpenStudy (anonymous):

yes, i'm just a little confused on the 5-(x+3) part

OpenStudy (anonymous):

so it would be is greater than or equal to 5 - -1x-3 which would be: -2x+6 is greater than or equal to -1x+8

OpenStudy (anonymous):

-2(x - 3) >= 5 -(x + 3) -2x + 6 >= 5 -x -3 -2x + 6 >= 2 - x -2x + x >= 2 - 6 -x >= -4 x <= 4

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

no. thank you : ) !

OpenStudy (anonymous):

that's exactly what I was going to say